# A generalization of the Choi isomorphism with application to open quantum systems

### v2

@@ I Introduction
After this breakthrough, it is not surprising that, as already alluded to in SMR61 , completely positive mappings are also used in the time evolution of open quantum systems. In 1976 two papers appeared simultaneously investigating semigroups of completely positive time evolutions, GKS76 and L76 , which led to the so-called Lindblad equation. This term has become commonplace, though perhaps not entirely fair. It is interesting for the history of the Lindblad equation that the first author of GKS76 , V. Gorini, became acquainted with the concept of completely positive mappings in discussions with K. Kraus at a conference in Marburg in 1973 HN74 . He told this to A. Kossakowski and informed E. Sudarshan, who invited to Austin both Kossakowski and Gorini for a few months in 1975-76, where the paper GKS76 was written VG24 . This part of the history of the Gorini-Kossakowski-Sudarshan-Lindblad equation is also touched in W25 .
For our purposes, eq. (2.4) in GKS76 is particularly interesting, since it establishes a 1:1 relationship between linear maps $\mathcal{E}$} of state spaces and Hermitean matrices, called GKS matrices $g$ in this paper, independent of the semigroup property. Consequently, the approach of GKS76 was later used to derive a time-dependent Lindblad equation which holds for general open quantum systems, see B12 or subsection VI.2 of this paper. We will prove that $\mathcal{E}$} is completely positive iff $g$ is positively semi-definite, thus generalizing Choi’s criterion, since the Choi matrix can be seen as a special case of the GKS matrix.
In quantum information theory, “quantum channels" are communication channels that can carry quantum information as well as classical information. Mathematically, they are defined as completely positive trace-preserving maps between spaces of operators, and thus as “operations" in the sense mentioned above. This definition can be traced back to H72 and has been therefore developed at the same time as the other physical applications mentioned above.
@@ IV.2 Definition of the GKS isomorphisms
${\mathcal{E}}\mapsto G({\mathcal{E}}):=g\;.$
@@ V Comparison with other isomorphisms
We will perform the same check with the GKS isomorphism corresponding to the orthonormal (Pauli) basis
in $\mathcal{A}$}. After some calculations we obtain the following GKS-matrix $G(\tau)$
@@ VI.2 Time-dependent Lindblad equation
${\mathcal{K}}(t)\rho(t)=-{\sf i}[H(t),\rho(t)]+\sum_{1\leq\alpha,\beta<N^{2}}{\sf k}^{\alpha\beta}(t)\left(F_{\alpha}\,\rho(t)\,F_{\beta}^{\ast}-\frac{1}{2}\left\{F_{\beta}^{\ast}F_{\alpha},\rho(t)\right\}\right)\;.$ (102)
@@ VII Summary
In the present work we have followed a different approach which can be traced back to the seminal paper GKS76 by Gorini, Kossakowski and Sudarshan, in which an early version of the Lindblad equation is derived. There an arbitrary superoperator $\mathcal{E}$} is written in the form
@@
In the second part of this work, we investigated consequences for the time evolution of general open quantum systems $S$ with environment $E$. Since this time evolution is completely positive, the corresponding time-dependent GKS matrix $g(t)$ must be positive semi-definite. This can also be confirmed directly by using a general Hamiltonian operator for the time evolution of the total system $S+E$, but only up to order $t^{2}$ due to the complexity of the calculations. The role of the GKS matrix for further physical applications of completely positive transformations, such as in the theory of the measurement process or quantum information theory, must be reserved for future investigations.

### v3

@@ I Introduction
After this breakthrough, it is not surprising that, as already alluded to in SMR61 , completely positive mappings are also used in the time evolution of open quantum systems. In 1976 two papers appeared simultaneously investigating semigroups of completely positive time evolutions, GKS76 and L76 , which led to the so-called Lindblad equation. This term has become commonplace, though perhaps not entirely fair. It is interesting for the history of the Lindblad equation that, acording to Vittorio Gorini VG24 , the first author of GKS76 became acquainted with the concept of completely positive mappings in discussions with K. Kraus at a conference in Marburg in 1973 HN74 . He told this to A. Kossakowski and informed E. Sudarshan, who invited to Austin both Kossakowski and Gorini for a few months in 1975-76, where the paper GKS76 was written. This part of the history of the Gorini-Kossakowski-Sudarshan-Lindblad equation is also touched in CP17 and in W25 .
For our purposes, eq. (2.4) in GKS76 is particularly interesting, since it establishes a 1:1 relationship between linear maps $\mathcal{E}$} of state spaces and Hermitean matrices, called GKS matrices $g$ in this paper, independent of the semigroup property. Consequently, the approach of GKS76 was later used to derive a time-dependent Lindblad equation which holds for general open quantum systems, see B12 or subsection VI.2 of this paper. We will prove that $\mathcal{E}$} is completely positive iff $g$ is positively semi-definite, thus generalizing Choi’s criterion, since the Choi matrix can be seen as a special case of the GKS matrix. This result is already partially included in papers on “‘quantum process tomography"; see eq. (6) in CN97 , eq. (8.152) in NC00 and Al03 , where the GKS matrix appears under the name “chi matrix."
In quantum information theory, “quantum channels" are communication channels that can carry quantum information as well as classical information. Mathematically, they are defined as completely positive trace-preserving maps between spaces of operators, and thus as “operations" in the sense mentioned above. This definition can be traced back to H72 and has been therefore developed at the same time as the other physical applications mentioned above. The Choi isomorphism is also known as “channel-state duality" in quantum information theory. It can be used, for example, to parameterize completely positive trace preserving maps for optimizing quantum processes, see HOK24 .
@@ IV.2 Definition of the GKS isomorphisms
${\mathcal{E}}\mapsto G({\mathcal{E}}):=g\;.$ (45)
@@
The above definition of the GKS-isomorphism is an implicit one. To obtain an explicit form, let us describe the superoperator ${\mathcal{E}}\in L({\mathcal{A}})$ by means of an $N^{2}\times N^{2}$-matrix $M$ in the usual way:
${\mathcal{E}}(F_{\alpha})=\sum_{\beta}M_{\beta\alpha}F_{\beta},\quad\mbox{for }0\leq\alpha<N^{2}\;.$ (46)
Further define a rank $4$ tensor $\Gamma$ by
$\Gamma^{\delta\gamma\sigma\tau}:=\mbox{Tr}\left(F_{\delta}F_{\sigma}^{\ast}F_{\gamma}^{\ast}F_{\tau}\right),\quad\mbox{for }0\leq\delta,\sigma,\gamma,\tau<N^{2}\;.$ (47)
Then, after some calculations, the implicit definition of the GKS-isomorphism can be transformed into the following
Definition 1
Let $\left(F_{\alpha}\right)_{\alpha=0,\ldots,N^{2}-1}$ denote an orthonormal basis in $\mathcal{A}$} w. r. t. the scalar product (1). Then the GKS-matrix isomorphism $M\mapsto g$ is defined by the equation
$g^{\delta\gamma}:=\sum_{\sigma\tau}\Gamma^{\delta\gamma\sigma\tau}\;M_{\sigma\tau},\quad\mbox{for }0\leq\delta,\gamma<N^{2}\;,$ (48)
with $\Gamma$ defined by (47) and $M$ by (46).
Sometimes the term “GKS matrix" appears in the literature with a slightly different meaning, namely in connection with the representation of the Lindblad generators of dynamic semigroups, see for example Betal01 . We will provide a note on this in the appropriate place, in subsection VI.2.
@@ IV.3 Properties of the GKS isomorphisms
As a simple example we consider the orthonormal (Pauli) basis
in $\mathcal{A}$} and the corresponding single qubit “Pauli channel"
${\mathcal{E}}(\rho)=\sum_{\alpha=0}^{3}k_{\alpha}F_{\alpha}\,\rho F_{\alpha}\;,$ (64)
where $k_{\alpha}>0$ for $\alpha=0,1,2,3$. It is trace-preserving iff $\sum_{\alpha=0}^{3}k_{\alpha}=2$. Obviously, the GKS-matrix of $\mathcal{E}$} is the diagonal matrix with entries $k_{0},k_{1},k_{2},k_{3}$, which is positive in accordance with the fact that the Pauli channel is completely positive.
@@ V Comparison with other isomorphisms
We will perform the same check with the GKS isomorphism corresponding to the orthonormal (Pauli) basis (63) in $\mathcal{A}$}. After some calculations we obtain the following GKS-matrix $G(\tau)$
@@ VI Application to the time evolution of open quantum systems
The time-independent GKSL equation for describing the dynamics of open quantum systems in the Markov approximation has numerous physical applications, which we cannot discuss here; for an overview, see, for example, AL07 . Instead, in this paper we will focus on the time-dependent GKSL equation and GKSL isomorphism.
@@ VI.2 Time-dependent Lindblad equation
${\mathcal{K}}(t)\rho(t)=-{\sf i}[H(t),\rho(t)]+\sum_{1\leq\alpha,\beta<N^{2}}{\sf k}^{\alpha\beta}(t)\left(F_{\alpha}\,\rho(t)\,F_{\beta}^{\ast}-\frac{1}{2}\left\{F_{\beta}^{\ast}F_{\alpha},\rho(t)\right\}\right)\;,$ (106)
using the anti-commutator $\{A,B\}:=AB+BA$ of Hermitean operators. If one makes a Markov approximation for the time evolution of the open system, then one can regard ${\mathcal{K}}(0)$ as the generator of a dynamic semigroup, cf. equation (2.3) in GKS76 or (7) in Betal01 . In this context, $k^{\alpha\beta}(0)$ is referred to as the “GKS matrix," see Betal01 , which is related to our notation but must nevertheless be distinguished.
@@ VII Summary
In the present work we have followed a different approach which can be traced back to the seminal paper GKS76 by Gorini, Kossakowski and Sudarshan, in which an early version of the Lindblad equation is derived, and has been re-discovered in the context of quantum process tomography CN97 , Al03 . There an arbitrary superoperator $\mathcal{E}$} is written in the form
@@
In the second part of this work, we investigated consequences for the time evolution of general open quantum systems $S$ with environment $E$. Since this time evolution is completely positive, the corresponding time-dependent GKS matrix $g(t)$ must be positive semi-definite. This can also be confirmed directly by using a general Hamiltonian operator for the time evolution of the total system $S+E$, but only up to order $t^{2}$ due to the complexity of the calculations. Investigations of the dynamics of open quantum systems beyond the Markov approximation are part of current research, see for example Setal25 . They are mostly carried out within the framework of the time-dependent Lindblad equation. However, there is a certain difference between the time-dependent Lindblad equation and the GKS equation for $g(t)$, although both equations are equivalent, as we have shown in subsection VI.2. The role of the GKS matrix for further physical applications of completely positive transformations, such as in the theory of the measurement process must be reserved for future investigations.